# Bridging the Synthesizability Gap in Perovskites by Combining Computations, Literature Data, and PU Learning

### v1

@@ 2.2 Positive Unlabeled Learning
To predict the synthesis likelihood of any material, one could utilize synthesis labels of known materials and derive similarity-based probability estimates. A label of “1” can be assigned to compounds with clear evidence of laboratory synthesis (as reported in the literature), but a label of “0” indicating clear non-synthesizability is not generally possible. For a given unknown compound, we can use the experimental positives and theoretical properties to estimate the likelihood of synthesis. To solve this task, in this work, we leveraged a semi-supervised machine learning technique called “Positive-Unlabeled Learning” or PU learning, which has been used in recent years by multiple researchers to predict synthesizability of a variety of solid-state materials [56, 59, 51, 60].
@@
As the name suggests, PU learning works upon a dataset with many positively labeled points and many unlabeled points. A classifier model is trained on the dataset by randomly assigning negative labels to unlabeled points. By learning from the combination of known and arbitrary labels, the model assigns an ensemble-based probability to the unlabeled points. There are many learning paradigms within this framework, and we used the “Transductive Bagging” approach introduced by Mordelet et al. [47]. In this approach, the first training loop considers all the known positive samples (K) and an equal number of random unlabeled samples (Un), which are assigned negative labels. The classifier learns the difference in their relations based on the input descriptors and assigns a probability to the remaining unlabeled points (U-Un) based on what has been learned. This step is repeated for T iterations, and the probability is then averaged over all the unlabeled points across the T loops. This workflow is pictured in Figure 4.
@@
Figure 4: Transductive Bagging for PU Learning. Each subset selection uses a K-fold split performed R times. For each K-fold in R, the model is trained for T iterations. The figure has been adapted from [52].
@@
As is evident from Figure 2(b), there is an imbalance in the DFT dataset in terms of compounds from different classes, since it is dominated by cubic ABX3 HaPs. To remove this bias while sampling the data, we used a repeated K-fold sampling approach during PU learning. In each repetition R, we applied K-fold sampling and T bagging iterations for each fold. This leads to better representation of each unlabeled point. The same approach is also applied when choosing the test set for the best classifier. This is necessary because a fixed test set, as was used in previous studies on synthesizability predictions [56, 51, 60], would introduce bias due to a class imbalance between types of perovskite compounds. ROC-AUC curves are eventually plotted to examine the training adequacy of the PU learning model. Since the positive label points are the only ones known with 100% reliability, only the “True Positive Rate” can be used as a valid metric on a sampled test set since all the unlabeled points are objectively unknown. We tested three different base-classifiers for training the PU learning models: Support Vector (SV), Random Forest (RF), and Decision Tree (DT).
@@
The positive labels applied to compounds in our dataset are obtained from two sources: ICSD tags from the MP database as described above, and labels collected from across the published literature. Scanning through potentially thousands of peer-reviewed publications to obtain relevant perovskite compositions is an extremely time-consuming task. Research on halide and chalcogenide perovskites is thriving, and new publications emerge every day. Tools based on natural language processing (NLP) [63, 64, 65, 66] have frequently been used for targeted data collection from textual information. While such tools are efficient at gathering particular information, such as chemical formulas and properties, it is still challenging to determine whether and how compounds were synthesized in the laboratory.
@@
State-of-the-art large language models (LLMs) are capable of efficiently extracting materials knowledge from large volumes of text [67]. Even so, processing entire PDFs of published papers using an LLM is time-consuming and exhausts many computational resources. Here, since our primary objective is to investigate the papers that report the synthesis of perovskite-type compositions, we only need to focus on the portions of text that list said compositions. This motivates our data collection pipeline pictured in Figure 6, where we extract chunks of text from a paper that contains a perovskite composition and pass them through an LLM. We used chain-of-thought inspired prompting, showing an example to the LLM to make it understand what we are looking for. For example, we directed the LLM to look for a composition that has an associated experimental method by giving it the names of the common characterization methods like X-ray diffraction (XRD), scanning electron microscope (SEM), etc.. The model is also shown keywords such as characterization patterns, deposition, experimental synthesis, etc., that can hint at the synthesis of a compound.
@@
Figure 6: Our workflow for extracting experimental labels from the literature using the Llama3.2 model [61] supported via the Ollama [68] library. Meaningful new compositions thus obtained are assigned a perovskite phase based on the literature or based on our predicted decomposition energy values.
@@
We first take the PDF of a publication and convert it into an XML format using docling [69], and then pass it through ChemDataExtractor2 [70], which uses Named Entity Recognition (NER) to filter portions that contain perovskite compositions. Here, the code tries to match the text to known chemical formulas of interest. We thus extract all the paragraphs that contain a perovskite-type chemical formula. The extracted data is then passed through an LLM using the Ollama [68] library that hosts multiple open-source local LLM schemes, providing a lightweight option to quickly screen through a large number of papers, and the output schema is handled at the end. If present, we extract the composition, crystal structure, evidence of synthesis, and properties of any compounds. Accurate collection of properties was found to be challenging and was discarded here in favor of simply obtaining synthesized perovskite compositions and phases. Using this pipeline, we processed over 1000 PDFs of perovskite literature collected by our group over the years and extracted more than 900 compositions.
@@
We found that there were many duplicates among the collected set of compounds, and sometimes the compositions were expressed as, e.g., “CsPbX3”, where X could be any halogen anion but was unspecified in the publication. It was also difficult to always parse the crystal structure information for the respective compounds, either due to model challenges or missing information. After cleaning and filtering the data, we obtained 206 unique compositions. Most of these compounds are alloys or contain different combinations of the previously discussed set of A/B/B’/X species. We kept the compositional space fixed when extracting the compounds so as not to obtain a sparsely represented dimension (ionic species). Since the perovskite descriptor needs both composition and phase, the final step was to determine the preferred crystal structure for the compositions with missing phase information. To do this, we used the DFT-computed or RGF-predicted decomposition energy of a compound and assigned the lowest energy phase as the ground state.
@@ 3 Results and Discussions
We tested three types of classifier models: a Support Vector (SV) classifier, a Random Forest (RF) classifier, and a Decision Tree (DT) classifier, each as implemented within the Scikit-learn library. To select the best base classifier, we rigorously trained three individual models and plotted their ROC curves along with the predicted probabilities for the test set. These models were trained on the combined dataset of unlabeled compounds (832 points) and compounds with experimental labels (239 points), within a probabilistic framework where (experimentally) unlabeled points were assigned random positive or negative labels through the ensemble learning process, as explained earlier. We found that each model performed significantly better than a purely random classifier, as pictured in Figure 7 (a), (b), (c). The DT classifier leads to the best performance with an area under the curve (AUC) of 0.908. This success can also be gleaned from the high true positive rate (TPR) of 86% for the DT model on the cumulative random test set, pictured in Figure 7 (f). In comparison, SV and RF models show TPR values of 81% and 84%, respectively. Thus, we used the DT classifier model going forward to make any new predictions. Optimum values of R = 10, K = 5, and T = 500 were found to yield the best AUC and TPR values for the DT model. Performance comparisons between different model parameters are provided in Figure S1.
We used the optimized DT classifier model to predict the synthesis probabilities for all the unlabeled samples, and the results are presented in Figure 8. 100 out of the 832 unlabeled compounds were predicted to have a synthesizability >> 0.5. A distribution of the different types of perovskites across these 100 compounds is shown in Figure 8(a), and a histogram of the predicted synthesizabilities across the dataset is pictured in Figure 8(b), showing that only a small fraction (12%) of the compounds have a greater than 50% likelihood of being made in the lab. Figure 8(c) shows the entire DFT dataset as a decomposition energy vs band gap plot like before, with the scatter points colored by their predicted synthesis probabilities.
@@
Table 1 shows the top 2 most synthesizable compounds predicted for each perovskite type (HaP, ChP, DP, VO). All compounds are cubic, have generally low decomposition energies, and photovoltaic-suitable band gaps with a couple of exceptions. There is a good variation in the types of cations observed in these compounds. There is a prevalence of Cs/FA and I/Br alloying among the most synthesizable HaP compounds, which has been shown in the past as a route to stabilize perovskites and reduce phase separation. For ChP ABX3 compounds, Sr is prevalent at the A site, and the compounds are generally sulfides or selenides. This is consistent with the recent thrust from researchers to focus on Sr-based perovskite chalcogenides [71] to move away from Ba-based compositions that suffer from larger than desired electron concentrations [72]. One of the vacancy-ordered double perovskites identified by the model is the compound Cs2PtBr6 [73], which is actually known to have been synthesized and was not included in our training dataset.
One of the main caveats about our synthesis prediction models is that it is almost definitely missing a notable number of compounds with positive labels that exist in the literature but have not been captured by our knowledge extraction framework, and adding all this data to the training set will certainly improve prediction accuracy and applicability. Furthermore, it is known that many halide and chalcogenide perovskites could exist in multiple phases at different conditions of temperature and pressure, while we currently only consider the single most stable phase at ambient conditions. It stands to reason that we could easily generate descriptors for compounds in multiple phases and obtain their synthesis probability values, which may reveal unexpected synthesizable phases for given compositions. Table 1 also lists some compounds such as Cs2ScAgBr6 and Cs2HfI6 which have not been previously synthesized (as far as we have seen in the literature) but are similar to other compositions that have indeed been made. A complete list of the 100 compounds with DFT properties that were predicted to have a synthesis probability of >> 50% is provided in the Supplementary Information.
@@
Finally, our regression models for DFT property prediction and the DT classifier model for synthesizability prediction are ready to be applied across the combinatorial space of all possible perovskite halides and chalcogenides within the defined chemical space of A/B/B’/X ions. To keep the problem tractable and for purposes of demonstration, we used only ABX3 compounds here. We enumerated a total of 20,000 novel compositions across different phases by applying constraints to the cation and anion choices to keep them within the chemical space we described above. We also ensure the compounds already existing in the dataset are not enumerated. Each composition enumerated is then repeated in the four different phases. Different cation and anion species were fit into the canonical ABX3 formula in appropriate oxidation states and mixing fractions to ensure overall charge neutrality. The generated compounds include unalloyed compositions as well as X-site alloys with only halogen anions, only chalcogen anions, and both halogen and chalcogen anions. The chalco-halide perovskites are included to demonstrate that our model can be utilized even for these cross-species materials, and it has been reported recently that chalco-halide compounds show desirable stability and defect tolerance [74]. Descriptors were generated for all 20,000 compounds, the RGF models were used to predict their (GGA-PBE accuracy) decomposition energy and band gap, and then the DFT properties were added to the descriptor list and the DT classifier model was used to predict their synthesis probabilities.
@@
Figure 9 shows histograms of the distribution of the predicted synthesizability across the space of the enumerated compounds, and a plot between the predicted decomposition energy and band gap of all these compounds, with scatter points colored by the predicted synthesis scores. We find that around 30% of the compounds have a synthesis probability >> 0.50, but the number of compounds drastically reduces for the highest probability values. Only 13% of the compounds show a value greater than 0.65. From Figure 9(b), it is clear that there is a correlation between the synthesis probability and the decomposition energy: nearly all the most synthesizable compounds have negative decomposition energies, and there is a noticable blue to red gradient from positive to negative decomposition energy values, similar to the original DFT dataset. This plot reveals hundreds of potential compositions with PV-suitable band gaps and high resistance to decomposition + synthesizability. Some of the top predicted compounds are highlighted in Table 2, with most of them adopting the cubic phase. Multiple compounds here are chalco-halides while the rest involve halogen mixing. The compound CsPbCl2Br has actually been reported experimentally before [75], thus providing some additional validation for our predictions. The predicted properties and synthesis probabilities of all 20,000 compounds are also provided in the SI.

### v2

@@ 2.2 Positive Unlabeled Learning
To predict the synthesis likelihood of any material, one could utilize synthesis labels of known materials and derive similarity-based probability estimates. A label of “1” can be assigned to compounds with clear evidence of laboratory synthesis (as reported in the literature), but a label of “0” indicating clear non-synthesizability is not generally possible. For a given unknown compound, we can use the experimental positives and theoretical properties to estimate the likelihood of synthesis. To solve this task, in this work, we leveraged a semi-supervised machine learning technique called “Positive-Unlabeled Learning” or PU learning, which has been used in recent years by multiple researchers to predict synthesizability of a variety of solid-state materials [pu-perovs, mof-pu, pucgcnn, pu-walsh].
@@
As the name suggests, PU learning works upon a dataset with many positively labeled points and many unlabeled points. A classifier model is trained on the dataset by randomly assigning negative labels to unlabeled points. By learning from the combination of known and arbitrary labels, the model assigns an ensemble-based probability to the unlabeled points. There are many learning paradigms within this framework, and we used the “Transductive Bagging” approach introduced by Mordelet et al. [pu-learning]. In this approach, the first training loop considers all the known positive samples (K) and an equal number of random unlabeled samples (Un), which are assigned negative labels. The classifier learns the difference in their relations based on the input descriptors and assigns a probability to the remaining unlabeled points (U-Un) based on what has been learned. This step is repeated for T iterations, and the probability is then averaged over all the unlabeled points across the T loops. This workflow is pictured in Figure 4.
@@
Figure 4: Transductive Bagging for PU Learning. Each subset selection uses a K-fold split performed R times. For each K-fold in R, the model is trained for T iterations. The figure has been adapted from [cgcnn].
@@
As is evident from Figure 2(b), there is an imbalance in the DFT dataset in terms of compounds from different classes, since it is dominated by cubic ABX3 HaPs. To remove this bias while sampling the data, we used a repeated K-fold sampling approach during PU learning. In each repetition R, we applied K-fold sampling and T bagging iterations for each fold. This leads to better representation of each unlabeled point. The same approach is also applied when choosing the test set for the best classifier. This is necessary because a fixed test set, as was used in previous studies on synthesizability predictions [pu-perovs, pucgcnn, pu-walsh], would introduce bias due to a class imbalance between types of perovskite compounds. ROC-AUC curves are eventually plotted to examine the training adequacy of the PU learning model. Since the positive label points are the only ones known with 100% reliability, only the “True Positive Rate” can be used as a valid metric on a sampled test set since all the unlabeled points are objectively unknown. We tested three different base-classifiers for training the PU learning models: Support Vector (SV), Random Forest (RF), and Decision Tree (DT).
@@
Elemental features such as electronegativity and boiling point provide the model with the chemical intuition for how the properties of constituent ions can affect the stability and synthesizability. On the other hand, the use of features such as ionic radii and the perovskite phase provides some structural information which helps assess phase stability better. The model could also benefit from specific synthesis conditions like temperature and pressure as well as information about likely synthesis routes and associated energy barriers, but the lack of consistent metadata about these aspects across the literature limits our ability to build a descriptor space from those quantities. Thus, the current descriptor choices only relate to the physical properties of the chemical species that make up any compound, and this would be sufficient for initial prediction and screening beyond which other synthetic information must be accounted for.
@@
The positive labels applied to compounds in our dataset are obtained from two sources: ICSD tags from the MP database as described above, and labels collected from across the published literature. The literature-based label extraction is crucial as updates on ICSD may lag behind publications for newly synthesized compositions, which can limit our positive labeled data. This is especially true for the general areas of halide and chalcogenide perovskites where the rate of new publications is substantial. Scanning through potentially thousands of peer-reviewed publications to obtain relevant perovskite compositions is an extremely time-consuming task. Research on halide and chalcogenide perovskites is thriving, and new publications emerge every day. Tools based on natural language processing (NLP) [chemdataextractor, nlp-1, nlp-2, nlp-3] have frequently been used for targeted data collection from textual information. While such tools are efficient at gathering particular information, such as chemical formulas and properties, it is still challenging to determine whether and how compounds were synthesized in the laboratory.
@@
State-of-the-art large language models (LLMs) are capable of efficiently extracting materials knowledge from large volumes of text [llm]. Even so, processing entire PDFs of published papers using an LLM is time-consuming and exhausts many computational resources. Here, since our primary objective is to investigate the papers that report the synthesis of perovskite-type compositions, we only need to focus on the portions of text that list said compositions. This motivates our data collection pipeline pictured in Figure 6, where we extract chunks of text from a paper that contains a perovskite composition and pass them through an LLM. We used chain-of-thought inspired prompting, showing an example to the LLM to make it understand what we are looking for. For example, we directed the LLM to look for a composition that has an associated experimental method by giving it the names of the common characterization methods like X-ray diffraction (XRD), scanning electron microscope (SEM), etc.. The model is also shown keywords such as characterization patterns, deposition, experimental synthesis, etc., that can hint at the synthesis of a compound.
@@
Figure 6: Our workflow for extracting experimental labels from the literature using the Llama3.2 model [Llama32] supported via the Ollama [Ollama] library. Meaningful new compositions thus obtained are assigned a perovskite phase based on the literature or based on our predicted decomposition energy values.
@@
We first take the PDF of a publication and convert it into an XML format using docling [Docling], and then pass it through ChemDataExtractor2 [chemdataextractorv2], which uses Named Entity Recognition (NER) to filter portions that contain perovskite compositions. Here, the code tries to match the text to known chemical formulas of interest. We thus extract all the paragraphs that contain a perovskite-type chemical formula. The extracted data is then passed through an LLM using the Ollama [Ollama] library that hosts multiple open-source local LLM schemes, providing a lightweight option to quickly screen through a large number of papers, and the output schema is handled at the end. If present, we extract the composition, crystal structure, evidence of synthesis, and properties of any compounds. Accurate collection of properties was found to be challenging and was discarded here in favor of simply obtaining synthesized perovskite compositions and phases. Using this pipeline, we processed over 1000 PDFs of perovskite literature collected by our group over the years and extracted more than 900 compositions.
@@
We found that there were many duplicates among the collected set of compounds, and sometimes the compositions were expressed as, e.g., “CsPbX3”, where X could be any halogen anion but was unspecified in the publication. It was also difficult to always parse the crystal structure information for the respective compounds, either due to model challenges or missing information. After cleaning and filtering the data, we obtained 206 unique compositions. Most of these compounds are alloys or contain different combinations of the previously discussed set of A/B/B’/X species. The relative distribution of cation and anion species in the positive and unlabeled sets is illustrated by the frequency charts in Figure S6, which reveal a similar distribution across both sets. This similarity ensures sufficient representation within the positive space, enabling reliable predictions of synthesizability for unlabeled compounds. We kept the compositional space fixed when extracting the compounds so as not to obtain a sparsely represented dimension (ionic species). Since the perovskite descriptor needs both composition and phase, the final step was to determine the preferred crystal structure for the compositions with missing phase information. To do this, we used the DFT-computed or RGF-predicted decomposition energy of a compound and assigned the lowest energy phase as the ground state.
@@ 3 Results and Discussions
We tested three types of classifier models: a Support Vector (SV) classifier, a Random Forest (RF) classifier, and a Decision Tree (DT) classifier, each as implemented within the Scikit-learn library. To select the best base classifier, we rigorously trained three individual models and plotted their ROC curves along with the predicted probabilities for the test set. These models were trained on the combined dataset of unlabeled compounds (832 points) and compounds with experimental labels (239 points), within a probabilistic framework where (experimentally) unlabeled points were assigned random positive or negative labels through the ensemble learning process, as explained earlier. We found that each model performed significantly better than a purely random classifier, as pictured in Figure 7 (a), (b), (c). The DT classifier leads to the best performance with an area under the curve (AUC) of 0.908. This success can also be gleaned from the high true positive rate (TPR) of 86% for the DT model on the cumulative random test set, pictured in Figure 7 (f). In comparison, SV and RF models show TPR values of 81% and 84%, respectively. Notably, some data points extracted from the literature populate low-probability regions. This is a useful outcome, as it indicates the model may be flagging false positives introduced by the automated extraction process. Thus, we used the DT classifier model going forward to make any new predictions. Optimum values of R = 10, K = 5, and T = 500 were found to yield the best AUC and TPR values for the DT model. Performance comparisons between different model parameters are provided in Figure S2.
We used the optimized DT classifier model to predict the synthesis probabilities for all the unlabeled samples, and the results are presented in Figure 8. 100 out of the 832 unlabeled compounds were predicted to have a synthesizability >> 0.5. A distribution of the different types of perovskites across these 100 compounds is shown in Figure 8(a), and a histogram of the predicted synthesizabilities across the dataset is pictured in Figure 8(b), showing that only a small fraction (12%) of the compounds have a greater than 50% likelihood of being made in the lab. Figure 8(c) shows the entire DFT dataset as a decomposition energy vs band gap plot like before, with the scatter points colored by their predicted synthesis probabilities. From the Figure 8(a), we can contend that our prediction spans across different perovskite types and crystal structures, even with limited positive labels, showing the robustness of this approach.
@@
Table 1 shows the top 2 most synthesizable compounds predicted for each perovskite type (HaP, ChP, DP, VO). All compounds are cubic, have generally low decomposition energies, and photovoltaic-suitable band gaps with a couple of exceptions. There is a good variation in the types of cations observed in these compounds. There is a prevalence of Cs/FA and I/Br alloying among the most synthesizable HaP compounds, which has been shown in the past as a route to stabilize perovskites and reduce phase separation. For ChP ABX3 compounds, Sr is prevalent at the A site, and the compounds are generally sulfides or selenides. This is consistent with the recent thrust from researchers to focus on Sr-based perovskite chalcogenides [sr-perovskites] to move away from Ba-based compositions that suffer from larger than desired electron concentrations [defect_bazrs3]. One of the vacancy-ordered double perovskites identified by the model is the compound Cs2PtBr6 [cs2ptbr6], which is actually known to have been synthesized and was not included in our training dataset.
One of the main caveats about our synthesis prediction models is that it is almost definitely missing a notable number of compounds with positive labels that exist in the literature but have not been captured by our knowledge extraction framework, and adding all this data to the training set will certainly improve prediction accuracy and applicability. Furthermore, it is known that many halide and chalcogenide perovskites could exist in multiple phases at different conditions of temperature and pressure, while we currently only consider the single most stable phase at ambient conditions. It stands to reason that we could easily generate descriptors for compounds in multiple phases and obtain their synthesis probability values, which may reveal unexpected synthesizable phases for given compositions. Table 1 also lists some compounds such as Cs2ScAgBr6 and Cs2HfI6 which have not been previously synthesized (as far as we have seen in the literature) but are similar to other compositions that have indeed been made. These compounds are already part of our dataset and have been optimized using GGA-PBE. Their relaxed structures are uploaded as part of SI and some of them are visualized in Figure S7. A complete list of the 100 compounds with DFT properties that were predicted to have a synthesis probability of >> 50% is provided in the Supplementary Information.
@@
Finally, our regression models for DFT property prediction and the DT classifier model for synthesizability prediction are ready to be applied across the combinatorial space of all possible perovskite halides and chalcogenides within the defined chemical space of A/B/B’/X ions. To keep the problem tractable and for purposes of demonstration, we used only ABX3 compounds here. We enumerated a total of 20,000 novel compositions across different phases by applying constraints to the cation and anion choices to keep them within the chemical space we described above. We also ensure the compounds already existing in the dataset are not enumerated. Each composition enumerated is then repeated in the four different phases. Different cation and anion species were fit into the canonical ABX3 formula in appropriate oxidation states and mixing fractions to ensure overall charge neutrality. The generated compounds include unalloyed compositions as well as X-site alloys with only halogen anions, only chalcogen anions, and both halogen and chalcogen anions. The chalco-halide perovskites are included to demonstrate that our model can be utilized even for these cross-species materials, and it has been reported recently that chalco-halide compounds show desirable stability and defect tolerance [chalcohalide]. Descriptors were generated for all 20,000 compounds, the RGF models were used to predict their (GGA-PBE accuracy) decomposition energy and band gap, and then the DFT properties were added to the descriptor list and the DT classifier model was used to predict their synthesis probabilities.
@@
Figure 9 shows histograms of the distribution of the predicted synthesizability across the space of the enumerated compounds, and a plot between the predicted decomposition energy and band gap of all these compounds, with scatter points colored by the predicted synthesis scores. We find that around 30% of the compounds have a synthesis probability >> 0.50, but the number of compounds drastically reduces for the highest probability values. Only 13% of the compounds show a value greater than 0.65. From Figure 9(b), it is clear that there is a correlation between the synthesis probability and the decomposition energy: nearly all the most synthesizable compounds have negative decomposition energies, and there is a noticable blue to red gradient from positive to negative decomposition energy values, similar to the original DFT dataset. This plot reveals hundreds of potential compositions with PV-suitable band gaps and high resistance to decomposition + synthesizability. Some of the top predicted compounds are highlighted in Table 2, with most of them adopting the cubic phase. Multiple compounds here are chalco-halides while the rest involve halogen mixing. The compound CsPbCl2Br has actually been reported experimentally before [cspbcl2br], thus providing some additional validation for our predictions. These models thus enable the study of various alloy compositions without the need for computationally expensive DFT calculations, allowing for the derivation of composition-property correlations. As an example, Figure S8 shows heatmaps of the RGF-predicted band gaps and predicted synthesis scores for Ba(Zr,Hf,Ti)S3 orthorhombic phase alloys, revealing some important information about how the incorporation of Zr/Hf/Ti at the B-site influences the stability and properties of interest. The predicted properties and synthesis probabilities of all 20,000 compounds are also provided in the SI.